As of July 2003, the population of India was estimated to be 1,050,000,000, with an annual growth rate of 1.47%. If the growth rate remains the same, how large will the population be in 50 years?
Answer by nyc_function(2741)

Use the standard compound interest formula.
A = P(1+r)^n
where A is the population
P is the principal or starting amount
r is the annual rate of growth
n is the number of years
A = 1050000000(1+0.0147)^50
A = 1050000000(1.0147)^50
A = 1050000000(2.074354695396870268818935813…
A = 2,178,072,430